OpenStudy (anonymous):

factor if possible: 4-81(x+y)^2

OpenStudy (anonymous):

(2-9(x+y))(2+9(x+y))

OpenStudy (anonymous):

4-81(x+y)^2 use formula a^2-b^2=(a-b)(a+b) (2)^2 - (9(x+y))^2 =(2+9(x+y))(2-9(x+y))